Assume that the wave is propagating from free space and striking to a dielectric medium normally. Give the derivation for reflection and transmission coefficient.


1
Expert's answer
2021-04-19T17:52:34-0400

Solution:-

To be asked in question

Reflection coefficient (R)=?

Transmission cofficient (T)=?

We know that fregenel equation

Reflection coefficient

"E_{or}=\frac{n_{1}-n_{2}}{n_{1}-n_{2}}E_{i}"

Transmission cofficient

"E_{oT}=\frac{n_{1}}{n_{1}+n_{2}}E_{i}"

Relation between c and v and n

"v=\frac{c}{n}"

Intensity

"I=\frac{1}{2}\epsilon vE^2"

We know that

R="\frac{I_{R}}{I_{I}}"

"R=[\frac{E_{OR}}{E_{oi}}]^2"

"R=(\frac {n_{1}-n_{2}}{{n_{1}+n_{2}}})^2"

Transmission cofficient

"T=\frac{I_{T}}{I_{R}}"

"T=\frac{\epsilon_{2}v_{2}}{\epsilon_{1}v_{1}}(\frac{E_0T}{E_0i})^2"

"T=\frac{4n_{1}n_{2}}{(n_{1}+n_{2})^2}"

"R+T=1"